Contents
What is the expected number of times that the coin lands on heads?
A coin is biased so that the probability of H is 0.6. What is the expected number of heads when it’s flipped 10 times? You can use an argument like that given in 2 using linearity of expectation, above, to see that the expectation is 10 times the expectation of H on one flip, that is, 10 times 0.6. So the answer is 6.
What is the expected number of heads that come up when a fair coin is flipped three times?
2265 answers so far. The probability of getting 3 heads when you toss a “fair” coin three times is (as others have said) 1 in 8, or 12. Let be the random variable denoting the number of heads obtained in 5 tosses. You flip a weighted coin that comes up H with probability 0.
What is the expected number of heads in four flips of a fair coin?
In counting the number of heads in 4 coin flips, the probability that we get exactly one head is the probability that we get anyone of the following 4 outcomes: HTTT, THTT, TTHT, or TTTH. Each has probability 1/16, so the probability to get exactly one head in 4 flips is 1/16 + 1/16 + 1/16 + 1/16 = 4/16 = 1/4.
What is the expected number of coin flips to get two heads?
The probability of this event is 1/4 and the total number of flips required will be 2. Framing the above three cases in the form of equations and adding we will get: Therefore, x = 6. Thus, the expected number of coin flips for getting two consecutive heads is 6.
What happens when the first coin flip has a tail?
In other words, the first tails makes all the previous tosses “wasted” and that increases the conditional expected time by that many tosses. Let the expected number of coin flips be . Now, there are three possible cases as listed below: If a tail appears on the first flip of coin.
How to calculate the probability of a coin flip?
We can condition E (X) on whatever our first flip is. Let E (X|H) denote the number of remaining coin flips given I got a head on the first flip. Similarly, let E (X|T) denote the number of remaining coin flips given I got a tail on the first flip.
How to calculate the expected number of coin tosses?
It uses a first time head case and a first time tail case to create a recursive equation and solves for E(x). E(x) = p + (1 − p)[1 + E(x)] where p is probability of getting heads, and E(x) is the expected number of coin tosses until landing on heads.